Filer Manufacturing has 9,893,589 shares of common stock outstanding. The current share price is $21.81, and the book value per share is $7.9. Filer Manufacturing also has two bond issues outstanding. The first bond issue has a face value of $67,593,513, has a 0.08 coupon, matures in 6 years and sells for 89 percent of par. The second issue has a face value of $41,607,683, has a 0.06 coupon, matures in 26 years, and sells for 92 percent of par. What is Filer's weight of equity on a market value basis? Enter the answer with 4 decimals (e.g. 0.2345)
Answer #1

weight of equity = Equity value/(total firm value)

total firm value=equity value+ bond 1 market value +bond 2 market value

=9,893,589 *21.81 +67,593,513*.89 + 41,607,683*.92 =314216471

weight of equity =9,893,589 *21.81/314216471 =.6867 =68.67%